### v2

@@ Abstract
If axion simultaneously has the scalar couplings to the nucleons and pseudo-scalar couplings to the electrons, it may mediate a $\mathcal{P}$, $\mathcal{T}$-violating interaction between the electronic shell and nuclei in the molecules. The polyatomic molecule RaOH, which is considered as a promising platform for the $\mathcal{P}$, $\mathcal{T}$ violation searches, is studied for its sensitivity to such interactions. Due to the long-range nature (on molecular scales) of the axion-mediated interaction, it is important whether the enhancement parameter would be sensitive to the vibration of the molecule. Our results imply that the impact of the vibrations on the axion-mediated electron-nucleon interaction in the molecule is similar to the impact on the short-range electron-nucleon scalar-pseudoscalar interaction studied earlier.
keywords:
axion , Dark matter , parity violation , time reversal symmetry violation , coupled channels method
[…]
[label1]organization=St. Petersburg State University,addressline=7/9 Universitetskaya nab., city=St. Petersburg, postcode=199034, state=, country=Russia
[…]
[label2]organization=Petersburg Nuclear Physics Institute named by B.P. Konstantinov of National Research Centre ”Kurchatov Institute”,addressline=1, mkr. Orlova roshcha, city=Gatchina, postcode=188300, country=Russia
1 Introduction
An important question in fundamental physics is the origin of Dark matter, the existence of which is predicted from the rotational curves of galaxies, and data from gravitational lensing of galaxy objects, such as the Bullet cluster. It also allows us to describe the formation of structures in the early Universe, and the observed inhomogeneities in cosmic microwave background radiation [1].
Pseudoscalar particles called axions are promising potential candidates for Dark matter constituents [2, 3, 4]. Such particles with certain masses and coupling constant values may solve the strong $\mathcal{C}$$\mathcal{P}$ problem [5], which remains an open problem of the Standard Model (SM) of particle physics [6, 7]. QCD does not have other sources of $\mathcal{C}$$\mathcal{P}$ violation except the axion-associated $\theta$-term. In the SM, all known parity violations occur in the weak interaction, related to the mixing of the quark matrices – Cabibbo-Kobayashi-Maskawa [8, 9], and the neutrino matrices – Pontecorvo-Maki-Nakagawa-Sakata [10, 11]. Some physical effects, such as the electron electric dipole moment (eEDM), and scalar-pseudoscalar electron-nucleon interaction (NE-SPS), are strongly suppressed in the Standard Model [12, 13, 14, 15]. However, some $\mathcal{CP}$-violating scenarios, such as SUSY, predict much larger values for these effects. The existence of new sources of $\mathcal{CP}$ violation may also be necessary to explain the problem of baryon asymmetry of the Universe, and may lead to a value of $\theta$-term closer to the neutron electric dipole moment limit [16].
@@ 1 Introduction
Molecular experiments allow us to study the effects related to $\mathcal{P}$ and $\mathcal{T}$-parity violation [17, 18, 13, 19, 20]. For example, an eEDM limit was obtained in an experiment with the diatomic molecule HfF+ [21]. Theoretical calculations are necessary to estimate the parameters of $\mathcal{P}$ and $\mathcal{T}$-parity violation for these effects [22]. The phenomena that may manifest themselves in the molecular spectra include the interaction with cosmic fields of axions and Dark photons, most straightforwardly in the chiral species such as the ground and vibrational states of the CHBrClF molecule [23].
@@
Among the studies of the effects caused by new interactions with axions in molecules, one can highlight the works on the effects of axion exchange between electrons and nuclei, as well as between electrons and electrons, studied for diatomic molecules HfF+ [24], BaF [25], and the triatomic molecule YbOH [26, 27]. Also, the interaction mediated by the axion produced by an electron-axion pseudoscalar vertex and a nucleon-axion or electron-axion scalar vertex was considered in [28, 29]. In these works, the contribution of such an interaction to the dipole moment of atoms and molecules was studied, and constraints on the corresponding interaction constants were obtained.
@@
It is important to stress that the measurement of the $\mathcal{P}$, $\mathcal{T}$ on the triatomic molecules involves their excited rovibrational states. In addition to the doublets associated with the rotation of the electron shell, triatomic linear molecules have $l$-doublets [30, 31, 32, 33]. In hexatomic molecules of the ”symmetric top” type, additional doublets appear associated with the rotation of the CH3 ligand in two opposite directions [34, 35]. Measurement of the relative shifts within different parity doublets allows one to suppress significantly many systematic effects. The possibility of laser cooling of the polyatomic molecules also presents a huge advantage for the experiment [36, 37, 38, 39, 40, 41].
@@
Because the states of interest are excited rovibrational states, it is important to study the impact of the molecule geometry on the enhancement parameters. Previously, we have studied such impact on the enhancement of the eEDM, and NE-SPS for the triatomic molecules RaOH [30], YbOH [31] as well as hexatomic molecules RaOCH3 [34] and YbOCH3[35]. While the impact is noticeable, it is not particularly large for these properties. The axion-mediated interaction differs from these interactions because of its long-range nature, therefore, one may find it interesting to check whether the impact of the molecular vibrations may become significant in this case. In this paper we study the enhancement parameter for the electron-nucleon axion-mediated interaction for the triatomic molecule RaOH.
2 Axion-mediated interaction of electrons and nuclei
@@
If both scalar interaction with nucleons $g_{N,S}$ (with $N=p,n$), and pseudo-scalar interaction with electrons $g_{e,P}$ are present, the process of axion exchange represented by a Feynman diagram in Fig. 1 induces the following one-electron operator of electron interaction with the nucleus $A$:
$\hat{H}_{eN}=i\frac{g_{e,P}}{4\pi}\sum_{i=1}^{N_{e}}\sum_{k=1}^{A}\int d^{3}\mathbf{R}_{k}\,g_{N_{k},S}\rho_{k}(\mathbf{R}_{k})\frac{e^{-m_{a}|\mathbf{r}_{i}-\mathbf{R}_{k}|}}{|\mathbf{r}_{i}-\mathbf{R}_{k}|}\gamma_{0}^{(i)}\gamma_{5}^{(i)},$ (2)
where $N_{e}$ is the number of electrons, $A$ - is the atomic weight of the nucleus (i.e. number of nucleons), $\mathbf{r}_{i}$ is the coordinate of the $i$-th electron, $\mathbf{R}_{k}$ is the coordinate of the $k$-th nucleon, $N_{k}=p,n$ is the type of nucleon, $\rho_{k}$ is the nucleon density function normalized by $1$, $m_{a}$ is the axion mass. The gamma matrices $\gamma_{0}^{(i)}$ and $\gamma_{5}^{(i)}$ are acting on the spinor of the $i$-th electron.
A number of approximations can be employed in this operator. First, the typical molecular length scale corresponds to the Compton length of masses in $\sim 1\,{\rm KeV}$ range. This is much larger than the astrophysically acceptable masses of the axion. Therefore, we can approximate the exponent in the numerator as $e^{-m_{a}|\mathbf{r_{i}}-\mathbf{R}_{k}|}\simeq 1$. Note that in the opposite limit of $m_{a}\gtrsim 10\,{\rm GeV}$ this interaction becomes an effective NE-SPS interaction [25].
We also approximate the nucleon density with a single function $\rho$, which we can identify as the charge density.
$\hat{H}_{eN}\simeq iA\frac{g_{e,P}\bar{g}_{N,S}}{4\pi}\sum_{i=1}^{N_{e}}\int d^{3}\mathbf{R}\,\frac{\rho(\mathbf{R})}{|\mathbf{r}_{i}-\mathbf{R}|}\gamma_{0}^{(i)}\gamma_{5}^{(i)},$ (3)
@@
In fact, the results of [25] imply that the non-point-like description of the atomic nuclei becomes important only for $m_{a}\gtrsim 10\,{\rm MeV}$. This is consistent with our own computation.
The enhancement parameter is then introduced as
$W_{\rm ax}=\frac{1}{g_{e,P}\bar{g}_{N,S}}\langle\Psi|\hat{H}_{eN}|\Psi\rangle,$ (5)
$\gamma_{5}$ presented in the operator requires electrons gaining relativistic momenta for this interaction to have any significance. Combined with $A$-scaling, this means that, despite its long-range nature, only the contribution of the interaction of electrons with heavy nuclei and their neighborhood (though significant in size) will matter. That’s why for the RaOH molecule, considered in this paper, we take into account only the contribution of the interaction of electrons with the heavy ($A=226$) Radium atom.
3 One-center restoration method
The computational complexity for the RaOH molecule is high due to the large number of electrons and the necessity to introduce large number of the basis functions to account for the states with large momenta. To overcome that, we use generalized effective core potential (GRECP) to describe the core electrons [42, 43, 44]. However, only the 2-component solutions can be obtained this way. And they have incorrect smoothed behavior in the core region. To overcome that problem, the one-center restoration method (OCR) was developed in [45]. Let us elucidate the main idea of this method.
@@ 3 One-center restoration method
The property is computed using the constructed 4-component molecular spinors.
4 Coupled-channels technique
@@
$q$ means coordinates of the electronic subsystem, $\hat{r}$ and $\hat{R}$ are directions of OH axis and Ra - center mass of OH axis respectively, $\theta$ is the angle between above axes, $R$ is the distance between the Radium atom and the ligand center of mass (Fig. 2). We consider the fixed ligand stretch since the Ra – OH bond stretching and bending frequencies are much smaller than the vibrations of the OH, using the equilibrium distance r=1.832a.u.r=1.832\,a.u. [46].
@@
Thus, we have a system of close-coupled equations for $F_{JjL}(R)$ [47]:
@@
where on the right side we have introduced a matrix potential,
@@
where a matrix with curly brackets denotes a $6j$-symbol, matrices with round brackets denote $3j$-symbols, and $V_{k}(R)$ correspond to the decomposition of the potential in terms of the Legendre polynomials $P_{k}(x)$,
@@
To deal with the dependence on $R$ we use the following approach. We choose a basis $f_{n}(R)$ of the eigenfunctions for longitudinal oscillations in the harmonic approximation,
@@
Then we use the following decomposition,
$F_{JjL}(R)=\sum_{n=0}^{n_{max}}F_{JjLn}f_{n}(R),$ (19)
and introduce matrix elements for the centrifugal term (from some small finite $R_{min}$ to large finite $R_{max}$) and potential,
@@
$\Bigg{[}-2\mu\omega_{\parallel}\Big{(}n+\frac{1}{2}\Big{)}-\frac{\mu j(j+1)}{\mu_{OH}r^{2}}+2\mu E\Big{]}F_{JjLn}\\ =2\sum_{n}L(L+1)C_{n\tilde{n}}F_{JjL\tilde{n}}+2\mu\sum_{\tilde{j},\tilde{l},\tilde{n}}\tilde{\mathcal{V}}_{jLn,\tilde{j}\tilde{L}\tilde{n}}^{J}F_{J\tilde{j}\tilde{L}\tilde{n}}$, (22)
@@
Applying similar processing (except 2nd step with Morse potential fit) and equation (23) to the property values, we also obtain the matrix elements of the properties $W_{jLn,\tilde{j}\tilde{L}\tilde{n}}$. The property average is obtained as,
@@
5 Results and discussion
The molecular orbitals were obtained in DIRAC19 software package using self-consistent field (SCF) level. To describe the heavy Radium atom we used a 10-valence electron basis set adapted to GRECP developed by the Quantum Physics and Chemistry Department of PNPI [48]. We have used the equivalent bases adapted to this potential previously used in [30].

### v3

@@ Abstract
If axion simultaneously has the scalar couplings to the nucleons and pseudo-scalar couplings to the electrons, it may mediate a $\mathcal{P}$, $\mathcal{T}$-violating interaction between the electronic shell and nuclei in the molecules. The polyatomic molecule RaOH, which is considered as a promising platform for the $\mathcal{P}$, $\mathcal{T}$-violation searches, is studied for its sensitivity to such interactions. Due to the long-range nature (on molecular scales) of the axion-mediated interaction, it is important whether the molecular parameter would be sensitive to the vibration of the molecule. Our results imply that the impact of the vibrations on the axion-mediated electron-nucleon interaction in the molecule is similar to the impact on the short-range electron-nucleon scalar-pseudoscalar interaction studied earlier.
I Introduction
An important question in fundamental physics is the origin of Dark matter, the existence of which is predicted from the rotational curves of galaxies, and data from gravitational lensing of galaxy objects, such as the Bullet cluster. It also allows us to describe the formation of structures in the early Universe, and the observed inhomogeneities in cosmic microwave background radiation Chadha-Day et al. (2022).
Pseudoscalar particles called axions are promising potential candidates for Dark matter constituents Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983). Such particles with certain masses and coupling constant values may solve the strong $\mathcal{C}$$\mathcal{P}$-problem Peccei and Quinn (1977), which remains an open problem of the Standard Model (SM) of particle physics Schwartz (2014); Navas et al. (2024). QCD does not have other sources of $\mathcal{C}$$\mathcal{P}$-violation except the axion-associated $\theta$-term. In the SM, all known parity violations occur in the weak interaction, related to the mixing of the quark matrices – Cabibbo-Kobayashi-Maskawa Cabibbo (1963); Kobayashi and Maskawa (1973), and the neutrino matrices – Pontecorvo-Maki-Nakagawa-Sakata Pontecorvo (1957); Maki et al. (1962). Some physical effects, such as the electron electric dipole moment (eEDM), and scalar-pseudoscalar electron-nucleon interaction (NE-SPS), are strongly suppressed in the Standard Model FUKUYAMA (2012); Pospelov and Ritz (2014); Yamaguchi and Yamanaka (2020, 2021). However, some $\mathcal{CP}$-violating scenarios, such as SUSY, predict much larger values for these effects. The existence of new sources of $\mathcal{CP}$-violation may also be necessary to explain the problem of baryon asymmetry of the Universe, and may lead to a value of $\theta$-term closer to the neutron electric dipole moment limit Sakharov (1998).
@@ I Introduction
Molecular experiments allow us to study the effects related to $\mathcal{P}$ and $\mathcal{T}$-parity violation Ginges and Flambaum (2004); Baron et al. (2014); Pospelov and Ritz (2014); Chubukov et al. (2019); Andreev et al. (2018). For example, an eEDM limit was obtained in an experiment with the diatomic molecule HfF+ Roussy et al. (2023). Theoretical calculations are necessary to estimate the parameters of $\mathcal{P}$ and $\mathcal{T}$-parity violation for these effects Kozlov and Labzowsky (1995). The phenomena that may manifest themselves in the molecular spectra include the interaction with cosmic fields of axions and Dark photons, most straightforwardly in the chiral species such as the ground and vibrational states of the CHBrClF molecule Gaul et al. (2020). Such interaction was also considered for a non-chiral molecule $RaOCH_{3}$in non-equilibrium configurations Zakharova (2025a)
@@
Among the studies of the effects caused by new interactions with axions in molecules, one can highlight the works on the effects of axion exchange between electrons and nuclei, as well as between electrons and electrons, studied for diatomic molecules HfF+ Prosnyak et al. (2023), BaF Prosnyak and Skripnikov (2024), and the triatomic molecule YbOH Maison et al. (2021a, b). Also, the interaction mediated by the axion produced by an electron-axion pseudoscalar vertex and a nucleon-axion or electron-axion scalar vertex was considered in Stadnik et al. (2018); Dzuba et al. (2018). In these works, the contribution of such an interaction to the dipole moment of atoms and molecules was studied, and constraints on the corresponding interaction constants were obtained.
@@
The polyatomic molecules such as the triatomic molecules RaOH and YbOH are a promising platform for the future parity violation searches. It is important to stress that the measurement of the $\mathcal{P}$, $\mathcal{T}$ on the triatomic molecules involves their excited rovibrational states. In addition to the doublets associated with the rotation of the electron shell, triatomic linear molecules have $l$-doublets Zakharova and Petrov (2021); Zakharova et al. (2021); Zakharova and Petrov (2022); Petrov and Zakharova (2022). In hexatomic molecules of the ”symmetric top” type, additional doublets appear associated with the rotation of the CH3 ligand in two opposite directions Zakharova (2022, 2024, 2025a). Measurement of the relative shifts within different parity doublets allows one to suppress significantly many systematic effects. The possibility of laser cooling of the polyatomic molecules also presents a huge advantage for the experiment Isaev and Berger (2016); Kozyryev et al. (2016, 2019); Augenbraun et al. (2021); Mitra et al. (2020); Fan et al. (2021). The laser cooling for the triatomic molecule SrOH was demonstrated in Sawaoka et al. (2025). For the molecule RaOH only limited spectroscopic data exists, such as in Conn et al. (2025) where the band origins, rotational constant and spin-rotation constant was obtained.
The axion-mediated interaction can manifest itself in the molecular spectra the same way as such P, T violating effects as the electron EDM and the scalar-pseudoscalar electron-nucleon interaction. It is important to compute the sensitivities to the different effects to be able to distinguish them from each other using multiple measurements on different molecular species.
If we suppose the existence of the axion exchange interaction of the electrons with nuclei characterized by the product of coupling constants $g_{e,P}\cdot g_{N,S}$, this $\mathcal{P}$, $\mathcal{T}$-odd effect can be estimated from the splitting between levels with opposite values of $M$ (projection of total angular momentum on space-fixed frame) given by:
$\Delta E_{\mathcal{P},\mathcal{T}}\simeq Pg_{e,P}g_{N,S}\Omega W_{\rm ax},$ (1)
@@
where $\Omega$ is the projection of the electron spin on the molecular axis. The parameter $W_{\rm ax}$ is determined from the molecular electronic structure calculation. $P\leq 1$ is the degree of polarization of the molecule Petrov and Zakharova (2022). Because the states of interest are excited rovibrational states, it is important to study the impact of the molecule geometry on the molecular parameters. Previously, we have studied such impact on the enhancement of the eEDM, and NE-SPS for the triatomic molecules RaOH Zakharova and Petrov (2021), YbOH Zakharova et al. (2021) as well as hexatomic molecules RaOCH3 Zakharova (2022) and YbOCH3Zakharova (2024). While the impact is noticeable, it is not particularly large for these properties. The axion-mediated interaction differs from these interactions because of its long-range nature, therefore, one may find it interesting to check whether the impact of the molecular vibrations may become significant in this case. In this paper we study the molecular parameter for the electron-nucleon axion-mediated interaction for the triatomic molecule RaOH.
II Axion-mediated interaction of electrons and nuclei
@@
If both scalar interaction with nucleons $g_{N,S}$ (with $N=p,n$), and pseudo-scalar interaction with electrons $g_{e,P}$ are present, the process of axion exchange represented by a Feynman diagram in Fig. 1 induces the following one-electron operator of electron interaction with the nucleus $A$, written in the atomic units as:
$\hat{H}_{eN}=i\frac{g_{e,P}}{4\pi}\sum_{i=1}^{N_{e}}\sum_{k=1}^{A}\int d^{3}\mathbf{R}_{k}\,g_{N_{k},S}\rho_{k}(\mathbf{R}_{k})\frac{e^{-m_{a}c|\mathbf{r}_{i}-\mathbf{R}_{k}|}}{|\mathbf{r}_{i}-\mathbf{R}_{k}|}\gamma_{0}^{(i)}\gamma_{5}^{(i)},$ (3)
where $N_{e}$ is the number of electrons, $A$ - is the atomic weight of the nucleus (i.e. number of nucleons), $\mathbf{r}_{i}$ is the coordinate of the $i$-th electron, $\mathbf{R}_{k}$ is the coordinate of the $k$-th nucleon, $N_{k}=p,n$ is the type of nucleon, $\rho_{k}$ is the nucleon density function normalized by $1$, $m_{a}$ is the axion mass. The gamma matrices $\gamma_{0}^{(i)}$ and $\gamma_{5}^{(i)}$ are acting on the spinor of the $i$-th electron. In the atomic units the speed of light is $c\simeq 137$.
A number of approximations can be employed in this operator.
First, we approximate the nucleon density with a single function $\rho$, which we can identify as the charge density.
$\hat{H}_{eN}\simeq iA\frac{g_{e,P}\bar{g}_{N,S}}{4\pi}\sum_{i=1}^{N_{e}}\int d^{3}\mathbf{R}\,\frac{\rho(\mathbf{R})e^{-m_{a}c|\mathbf{r}_{i}-\mathbf{R}_{k}|}}{|\mathbf{r}_{i}-\mathbf{R}|}\gamma_{0}^{(i)}\gamma_{5}^{(i)},$ (4)
@@
In fact, the results of Prosnyak and Skripnikov (2024) imply that the non-point-like description of the atomic nuclei becomes important only for $m_{a}\gtrsim 10\,{\rm MeV}$. This is consistent with our own computation.
The molecular parameter is then introduced as
$W_{\rm ax}=\frac{1}{g_{e,P}\bar{g}_{N,S}{\Omega}}\langle\Psi|\hat{H}_{eN}|\Psi\rangle,$ (6)
$\gamma_{5}$ presented in the operator requires electrons gaining relativistic momenta for this interaction to have any significance. Combined with $A$-scaling, this means that, despite its long-range nature, only the contribution of the interaction of electrons with heavy nuclei and their neighborhood (though significant in size) will matter. That’s why for the RaOH molecule, considered in this paper, we take into account only the contribution of the interaction of electrons with the heavy ($A=226$) Radium atom. The Eq.(5) gives the expectation value of the axion- mediated $\mathcal{P}$ and $\mathcal{T}$-violating part of Hamiltonian. This value can be used to obtain product of constants $g_{e,P}\cdot g_{N,S}$ from energy shifts of molecular levels that could, in principle, be detected experimentally.
Second, the typical molecular length scale corresponds to the Compton length of masses in $\sim 1\,{\rm KeV}$ range. This is much larger than the astrophysically acceptable masses of the axion. Therefore, we can approximate the exponent in the numerator as $e^{-m_{a}|\mathbf{r_{i}}-\mathbf{R}_{k}|}\simeq 1$. However, it is still interesting to consider the opposite limit of $m_{a}c\gtrsim 10\,{\rm GeV}$ in which this interaction becomes an effective NE-SPS interaction Prosnyak and Skripnikov (2024):
$\hat{H}_{s}=ig_{s}\frac{G_{F}}{\sqrt{2}}Z\rho(\vec{r})\gamma_{0}\gamma_{5}$ (7)
with the parameter describing the sensitivity to this interaction:
$W_{s}=\frac{1}{g_{s}\Omega}\langle\Psi|\hat{H}_{s}|\Psi\rangle,$ (8)
where the Fermi constant is $G_{F}=2.2225\cdot 10^{-14}$ in the atomic units. In the high-mass limit when $m_{a}c$ comparable and less to the radius of the nucleus, these parameters are related to each other as
$W_{\rm ax}\simeq\frac{\sqrt{2}A}{ZG_{F}}\frac{W_{\rm s}}{m_{a}^{2}c^{2}}$ (9)
As we have previously computed the value of $W_{s}$ in Zakharova and Petrov (2021), we will use this relation to check the validity of our computation.
III One-center restoration method
The computational complexity for the RaOH molecule is high due to the large number of electrons and the necessity to introduce large number of the basis functions to account for the states with large momenta. To overcome that, we use generalized effective core potential (GRECP) to describe the core electrons Titov and Mosyagin (1999); Mosyagin et al. (2010, 2016). However, only the 2-component solutions can be obtained this way. And they have incorrect smoothed behavior in the core region. To overcome that problem, the one-center restoration method (OCR) was developed in Titov et al. (2006). Let us elucidate the main idea of this method.
@@ III One-center restoration method
The property is computed using the constructed 4-component molecular spinors. Details of implementation are given in Zakharova (2025b).
IV Coupled-channels technique
@@
$q$ means coordinates of the electronic subsystem, $\hat{r}$ and $\hat{R}$ are directions of OH axis and Ra - center mass of OH axis respectively, $\theta$ is the angle between above axes, $R$ is the distance between the Radium atom and the ligand center of mass (Fig. 2). We consider the fixed ligand stretch since the Ra – OH bond stretching and bending frequencies are much smaller than the vibrations of the OH, using the equilibrium distance r=1.832a.u.r=1.832\,a.u. Huber and Herzberg (1979).
@@
Thus, we have a system of close-coupled equations for $F_{JjL}(R)$ McGuire and Kouri (1974):
@@
where on the right side we have introduced a matrix potential:
@@
where a matrix with curly brackets denotes a $6j$-symbol, matrices with round brackets denote $3j$-symbols, and $V_{k}(R)$ correspond to the decomposition of the potential in terms of the Legendre polynomials $P_{k}(x)$:
@@
To deal with the dependence on $R$ we use the following approach. We choose a basis $f_{n}(R)$ of the eigenfunctions for longitudinal oscillations in the harmonic approximation:
@@
Then we use the following decomposition:
$F_{JjL}(R)=\sum_{n=0}^{n_{max}}F_{JjLn}f_{n}(R).$ (23)
The exact solution would correspond to the infinite sum but for the numerical computations we use the finite basis approximation, and cut our sum to $n_{max}$. We considered different values of $n_{max}$ to check numerical convergence. Here $n_{max}$ is a maximal occupation number for the longitudinal vibrations. This means that higher value of $n_{max}$ describes wider wavefunctions as well as more rapid oscillations in $R$. The electronic computations are performed on a discrete grid of $R$ with minimal and maximal values defined $R_{min}$ and $R_{max}$ correspondingly. We introduce matrix elements for the centrifugal term is:
@@
The values for $R_{min}$ and higher $R_{max}$ are chosen in such a way that the obtained nuclear motion wavefunctions decrease rapidly beyond that range and do not impact the accuracy of the numerical computation.
@@
$\Bigg[-2\mu\omega_{\parallel}\Big(n+\frac{1}{2}\Big)-\frac{\mu j(j+1)}{\mu_{OH}r^{2}}+2\mu E\Big]F_{JjLn}=2\sum_{n}L(L+1)C_{n\tilde{n}}F_{JjL\tilde{n}}+2\mu\sum_{\tilde{j},\tilde{l},\tilde{n}}\tilde{\mathcal{V}}_{jLn,\tilde{j}\tilde{L}\tilde{n}}^{J}F_{J\tilde{j}\tilde{L}\tilde{n}}$, (26)
@@
Applying similar processing (except second step with Morse potential fit) and equation (27) to the property values, we also obtain the matrix elements of the properties $W_{jLn,\tilde{j}\tilde{L}\tilde{n}}$. The property average is obtained as
@@
V Results and discussion
The molecular orbitals were obtained in DIRAC19 software package using self-consistent field (SCF) level. To describe the heavy Radium atom we used a 10-valence electron basis set adapted to GRECP developed by the Quantum Physics and Chemistry Department of PNPI URL: http://www.qchem.pnpi.spb.ru/Basis/ . We have used the equivalent bases adapted to this potential previously used in Zakharova and Petrov (2021).